OpenStudy (anonymous):

A longer side of a parallelogram is 4dm shorter than 5 times the length of a shorter side. One side of a Rhombus (Equilateral Quadrilateral) and a longer side of the parallelogram are equal in length. If the difference in the two perimeters is 48dm, find the perimeter of the rhombus.

OpenStudy (azureilai):

P=2l+2w l=4-5w P1-P2=48 You can use above equations to solve